OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 5n + 6 = -4

OpenStudy (anonymous):

5n+6=-4 5n+6-6=-4-6 5n=-10 5n/5=-10/5 n=-2

OpenStudy (anonymous):